solve the following systems for the values of x and y:
y-x=1
y=1/2x+4

Answer by checkley79(3341)

y-x=1 OR y=x-1
y=1/2x+4
x-1=x/2+4
X-X/2=4+1
X/2=5
X=5*2
X=10 ANS.
Y=10-1
Y=9 ANS.
PROOF:
9=10/2+4
9=5+4
9=9